# Persistence of AR($1$) sequences with Rademacher innovations and linear mod $1$ transforms

### v2

@@ 1.2. Main results
    For $a=\frac{2}{3}$, the mapping $T_{2/3}$ is defined on the whole of the interval $[0,\frac{1}{1-a}]$. In this case we can consider our problem from a different perspective, since the killed transition operator $P$ turns out to be a weighted transfer (or Ruelle) operator. For $p=\frac{1}{2}$, we can specify that $P=\frac{3}{4}P_{T}$, where $P_{T}$ is the transfer (or Perron–Frobenius) operator of $T=T_{2/3}$. For this $T$, it is well known that $P_{T}$ is quasi-compact on the space of functions of bounded variation on $[0,3]$; see e.g. Boyarsky and Góra [6]. The leading eigenvalue of $P_{T}$ is simple and equals $1$, and the corresponding eigenfunction is the invariant density of the transformation $\hat{T}_{2/3}$ rescaled to the domain $[0,3]$. This density is explicitly given (up to a multiplicative factor) by formula (58), which is due to Parry [19]. We recover $V$ by rescaling and simplifying (58); notably, in this case $cV$ is a probability density. We also get $\lambda_{2/3}(1/2)=3/4$ by $P=(3/4)P_{T}$. Lastly, the quasi-invariant distribution $\nu_{2/3}$ is uniform when $p=1/2$ because the uniform density is an eigenfunction of the composition (or Koopman) operator that is dual to $P_{T}$; there are no other eigenfunctions by [6, Theorem 3.5.2].
  5. (e)
    The constant $c$ is given in (31).
  6. (f)
    The value of $\lambda_{a}$ is rather explicit from (11) when $\varkappa_{a}^{\prime}<\infty$, and we will give concrete examples in Section 2.3. In this case the l.h.s. in (11) can be written as a finite sum even if $\varkappa_{a}=\infty$ because then the sequence $\{\delta_{k}\}$ is eventually periodic. If $\varkappa_{a}^{\prime}=\infty$, no simplification of (11) appears to be possible because of chaotic behaviour of the orbit of zero unless $a=\frac{2}{3}$ and $p=\frac{1}{2}$. In this case we found $\lambda_{a}$ explicitly, computing the l.h.s. in (11) when $\lambda=ap$ using the $(\frac{1}{a},\frac{1}{2})$-expansion of 00. This gives $\lambda_{2/3}(1/2)=3/4$.
  7. (g)
    The rate of convergence in (10) and  (13) is geometric and uniform in $x$ and $y$; see (63). Moreover, (13) extends to convergence of $\mathbb{E}_{x}(f(X_{n})|\tau>n)$, where the rate of convergence is uniform in $x$ on sets of functions $f$ of uniformly bounded variation; see Proposition 16. We give a full proof of this result only for $p\geq 1/2$.
  8. (h)
    As we have already mentioned, if the starting point $x$ exceeds $1/(1-a)$, the chain strictly decreases until the stopping time
    $\sigma=\inf\{n\geq 0:X_{n}\leq 1/(1-a)\}.$
    This random variable is finite a.s. because $\sigma$ is the index of the first negative innovation after the chain downcrosses the higher level $(2-a)/(a(1-a))$. This downcrossing happens in a bounded number of steps (which depends on the starting point). Moreover, one can easily show that for some constant $C(x)>0$,
    $\mathbb{P}_{x}(\sigma>k)\leq C(x)p^{k}.$
    Using first the strong Markov property, and then applying (10) and the above bound for the tail probability of $\sigma$, we conclude that for any $a\in(\frac{1}{2},\frac{2}{3}]$,
    $\mathbb{P}_{x}(\tau>n)\sim c\lambda_{a}^{n}\sum_{k=0}^{\varkappa_{a}}\Big{(}\frac{p}{\lambda_{a}}\Big{)}^{k}\Big{(}\frac{q}{p}\Big{)}^{L_{k}}\mathbb{E}_{x}[\lambda_{a}^{-\sigma}\mathbbm{1}\{T_{a}^{k}(0)\leq X_{\sigma}\}].$
  9. (i)
@@
    For $a\in(0,\frac{1}{2}]$, the analysis of persistence is rather simple. Indeed, in this case $\frac{1}{1-a}\leq\frac{1}{a}$, hence for any starting point $x\in[0,\frac{1}{a})$ it holds that $\tau=\inf\{n\geq 1:\xi_{n}=-1\}$. Consequently, for such $x$ we have $\mathbb{P}_{x}(\tau>n)=p^{n}$, and thus $\lambda_{a}(p)=p$ for $a\in(0,\frac{1}{2}]$.
    To find the tail asymptotics for $x\geq 1/a$, define $\sigma^{\prime}=\inf\{n\geq 0:X_{n}<1/a\}$. This random variable is bounded when $a<\frac{1}{2}$ because $X_{n}<xa^{n}+\frac{1}{1-a}$ and in this case $\frac{1}{1-a}<\frac{1}{a}$. Therefore, for all $x\geq 0$ when $a<\frac{1}{2}$ and for all $20\leq x<$ 2 when $a=\frac{1}{2}$,
    $\mathbb{P}_{x}(\tau>n)=\mathbb{E}_{x}p^{-\sigma^{\prime}}\cdot p^{n}.$
    For $a=\frac{1}{2}$ and $x\geq 2$, the variable $\sigma^{\prime}$ is not bounded and $\mathbb{E}_{x}p^{-\sigma^{\prime}}=\infty$. However, we can write $\sigma^{\prime}=\inf\{n>\sigma^{\prime\prime}:\xi_{n}=-1\}$, where $\sigma^{\prime\prime}=\inf\{n\geq 0:2\leq X_{n}<6\}$ is bounded. Hence, invoking the tail asymptotics of the sum of independent geometric random variables $\tau-\sigma^{\prime}$ and $\sigma^{\prime}-\sigma^{\prime\prime}$, it is easy to obtain that
    as $n\to\infty$. Notably, the order of this asymptotics differs from that in (10).
    It is easy to see that if $a<1/2$, then the quasi-stationary distribution $\nu_{a}$ is the delta-measure at point $\frac{1}{1-a}$.
  10. (j)
@@
    Our approach can be extended to cover persistence of the AR(1) chains with i.i.d. innovations that take two arbitrary values. Indeed, w.l.o.g. assume that $\mathbb{P}(\xi_{1}=1)=p$ and $\mathbb{P}(\xi_{1}=-b)=q$ for some $b>0$. Then for any $a\in(\frac{b}{b+1},\frac{b+1}{b+2})$, the deterministic relation in reversed time (6) remains valid if we substitute $T_{a}$ by the mapping $T_{a,b}$ that is defined by
@@
    It is not hard to check that for such $a$, relations (10), (11), (12) remain valid with $T_{a}$ and $\varkappa_{a}$ substituted by $T_{a,b}$ and $\varkappa_{a,b}$ if we define $\varkappa_{a,b}$ as in (7) with $I_{a}$ replaced by $(\frac{a(b+1)-b}{1-a},1)$. In contrast, the study of $\lambda_{a}$ reduces to analysis of the mapping $\hat{T}_{\beta,\alpha}$ with $\beta=1/a$ and $\alpha=b/(b+1)$, whose properties substantially depend on $b$. $\diamond$
The structure of the paper is clear from the titles of the following sections.
2. Explanation of our methods and related discussions
In this section we give three different ways to introduce the mapping $T_{a}$ and explain its role in our solution of the exit problem for the AR($1$) chain with ±$\pm 1$ innovations.
@@ 2. Explanation of our methods and related discussions
2.1. Deterministic behaviour in reversed time
@@
Let us prove (6) for $a\in[\frac{1}{2},\frac{2}{3})$. If $X_{n}<1$, then $X_{n}=aX_{n-1}-1$ because the other option $X_{n}=aX_{n-1}+1$ is impossible by $aX_{n-1}+1\geq 1$ on $\{\tau>n\}$. Hence $X_{n}<1$ by (14). Thus, $X_{n}$ is in the domain of $T_{a}$ and we can write $X_{n-1}=\frac{1}{a}(X_{n}+1)=T_{a}(X_{n})$. If $X_{n}\geq 1$, then $X_{n}=aX_{n-1}+1$ because otherwise $X_{n}=aX_{n-1}-1<1$ by (14). Then $X_{n-1}=\frac{1}{a}(X_{n}-1)=T_{a}(X_{n})$. We thus checked that $X_{n-1}=T_{a}(X_{n})$ holds true on $\{\tau>n\}$ in either case, and (6) follows by induction.
@@
$\xi_{n-k}=(-1)^{1-d_{k}(aX_{n}/2)}=2d_{k}(aX_{n}/2)-1.$
Thus, $(\xi_{n-k}+1)/2$ for $10\leq k\leq n-$ 1 are the first $n$ digits in the $(\frac{1}{a},\frac{1}{2})$-expansion of $aX_{n}/2$.
The deterministic dynamics in the reversed time described above is essentially the reason for singular continuity of the quasi-stationary distributions $\nu_{a}$. The same effect can be also observed in the unconditional setting. More precisely, we shall now show that $X_{n}$ is deterministic in the reversed time for every $a\in(0,\frac{1}{2})$. We first notice that $|X_{k}|\geq\frac{1-2a}{1-a}$ for all $k\geq 1$ and for all starting points $x\in[-\frac{1}{1-a},\frac{1}{1-a}]$; this follows as in (14). Assume that $X_{n}\geq\frac{1-2a}{1-a}$. In general, we have two possibilities for $X_{n-1}$, namely $X_{n-1}=\frac{1}{a}(X_{n}+1)$ and $X_{n-1}=\frac{1}{a}(X_{n}-1)$. The assumption $X_{n}\geq\frac{1-2a}{1-a}$ implies that the case where $X_{n-1}=\frac{1}{a}(X_{n}+1)$ is not possible since
@@
This property implies that the limiting distribution $\pi_{a}$ is supported on the points $z$ whose orbit under $G_{a}$ does not hit the interval $\left(-\frac{1-2a}{1-a},\frac{1-2a}{1-a}\right)$. The set of such points has Lebesgue measure zero. Since the two “branches” of $G_{a}$ are onto the whole of the interval $[-\frac{1}{1-a},\frac{1}{1-a}]$, besides showing that the distribution $\pi_{a}$ is singular and continuous, one can compute the Hausdorff dimension of its support. The proof of these properties via deterministic evolution in the reversed time can be found in Lemma 2 by Bovier [8]. Relation (II-17) there is equivalent to (15).
2.2. The compactness approach
Let $P$ be the transition operator of $X_{n}$ killed at exiting $[0,\frac{1}{1-a}]$. It acts on measurable bounded functions on $[0,\frac{1}{1-a}]$ as follows:
$Pf(x)=\mathbb{E}_{x}[f(X_{1});\tau>1].$
Under assumption (4) that the innovations are ±$\pm 1$, we have
$Pf(x)=pf(ax+1)+qf(ax-1)\mathbbm{1}\{x\geq 1/a\}.$ (16)
Since
$\mathbb{P}_{x}(\tau>n)=P^{n}\mathbbm{1}(x),$
we seek to find a Banach space of functions on $[0,1/(1-a)]$ where $P$ would be compact or quasi-compact, expecting that the asymptotics of $P^{n}\mathbbm{1}$ is defined by the spectral radius of $P$, which should be the largest eigenvalue of $P$. To this end, let us consider the behaviour of the set of discontinuities $D_{f}$ of a function $f$ under the action of $P$, with continuity at the endpoints 00 and $1/(1-a)$ understood as one-sided continuity.
1. The case $a\in(0,\frac{2}{3})$.
Assume that $a\in[\frac{1}{2},\frac{2}{3})$. If $Pf$ is discontinuous at an $x\in[0,\frac{1}{1-a}]$, then the same holds true for at least one of the two terms in (16). Note that $f(ax+1)$ is discontinuous at $x$ if and only if $f$ is discontinuous at $y=ax+1$; here $y\in[1,\frac{1}{1-a}]$ and $x=\frac{1}{a}(y-1)$. On the other hand, $f(ax-1)\mathbbm{1}\{x\geq 1/a\}$ is discontinuous at $x$ if and only if $f$ is discontinuous at $y=ax-1$ or it holds that $x=1/a,f(0)\neq 0$; here $y\in[0,\frac{2a-1}{1-a}]$ and $x=\frac{1}{a}(y+1)$ in both cases.
The observations above can be summed up using the mapping
$D_{Pf}\subset T_{a}\big{(}(D_{f}\cup\{0\})\setminus I_{a}\big{)}.$ (17)
It follows from (17) that for $a\in(0,\frac{2}{3})$, the set of measurable bounded functions on $[0,\frac{1}{1-a}]$ that are continuous at every point outside of $\{T_{a}^{k}(0):0\leq k<\varkappa_{a}+1\}$ is closed under the action of $P$; recall that
$\varkappa_{a}=\inf\{k\geq 0:\,T_{a}^{k}(0)\in I_{a}\}.$
This suggests to consider the set of functions
on $[0,\frac{1}{1-a}]$, because we know how $P$ transforms the jumps of functions and every function in $U$ is defined by its jumps. The idea to consider general right-continuous step functions with countably many jumps (called saltus functions) in the context of linear mod 1 transforms goes back to Halfin [15].
We will show that the set $U$ is closed under the action of $P$. We will give a simple explicit description of this action and show that $P$ is a quasi-compact operator on $U$.
2. The case $a\in[\frac{2}{3},1)$.
Here $\frac{2a-1}{1-a}\geq 1$ and thus the mapping $T_{a}$ does not account for the discontinuities of $Pf$ arising from the discontinuities of $f$ on $[1,\frac{2a-1}{1-a}]$ due to the second term in (16). The argument above gives
$D_{Pf}\subset T_{a}\Big{(}D_{f}\cup\{0\}\Big{)}\cup\Big{\{}\frac{1}{a}(y+1):y\in D_{f}\cap\Big{[}1,\frac{2a-1}{1-a}\Big{]}\Big{\}}.$ (19)
The r.h.s. simplifies when $f$ has no discontinuities in $[1,\frac{2a-1}{1-a}]$. This motivates us to introduce
$\varkappa_{a}=\inf\Big{\{}k\geq 0:\,T_{a}^{k}(0)\in\Big{[}1,\frac{2a-1}{1-a}\Big{]}\Big{\}}.$
which holds true for every $a\in[\frac{1}{2},\frac{2}{3}]$ and follows by simple induction. Hence if $T_{2/3}^{k}(0)=1$, then $2^{k}=-\sum_{i=0}^{k-1}(-1)^{\delta_{i}}3^{k-i}2^{i}$, which is impossible because the r.h.s. is odd.
With this extended definition of $\varkappa_{a}$, the functional space $U$ introduced in (18) is again closed under the action of $P$ if $a\in[\frac{2}{3},1)$ is such that $\varkappa_{a}=\infty$. The set of such $a$ is contained in $[\frac{2}{3},\frac{\sqrt{2}}{2})$ because $T_{a}(0)\not\in[1,\frac{2a-1}{1-a}]$ only when $\frac{2a-1}{1-a}<\frac{1}{a}$, hence $12a^{2}<$ 1. It is possible to show that this set has Lebesgue measure zero. Note that it contains points other than $a=2/3$, for example, the unique solution to $T_{a}^{5}(0)=1/a$ on $[2/3,1)$, which is $a\approx 0.691$.
Remark 3.
Our method of proving tail asymptotics (10) seem to work unchanged for every $a\in(\frac{2}{3},1)$ such that $\varkappa_{a}=\infty$. $\diamond$
Remark 4.
The r.h.s. reminds of a weighted Perron–Frobenuis operator, where the weight takes two values $p$ and $q$; such operators are considered e.g. in the book by Baladi [4]. However, the actual operator $P$ is defective because the range of the mapping $T_{a}$ is larger than its domain, which has a “hole” $I_{a}$. This hole vanishes when $a=\frac{2}{3}$. In this case we can consider the standard Perron–Frobenius operator $P_{T_{a}}$ for $T_{a}$, given by $P_{T_{a}}f(x)=\sum_{y\in T_{a}^{-1}(x)}af(y)$. This is the density of the measure with density $f$ transformed (pushed forward) by $T_{a}$. Then for $a=\frac{2}{3}$ and $p=\frac{1}{2}$ we have
$Pf(x)=\frac{3}{4}P_{T_{a}}f(x),\qquad 0\leq x<3.$
2.3. The aggregated chain approach

### v3

@@ 1.2. Main results
    The rate of convergence in (14) and  (17) is exponential; see (43) and (75). Moreover, we extend (17) to convergence of the conditional functionals of the form $\mathbb{E}_{x}(f(X_{n})|\tau>n)$, which also holds true at an exponential rate; see Proposition 18.
@@
We now consider the case $a\in[0,\frac{1}{2}]$. Here, the analysis of persistence is rather straightforward. By $\frac{1}{1-a}\leq\frac{1}{a}$, for any starting point $x\in[0,\frac{1}{a})$, we have $\tau=\inf\{n\geq 1:\xi_{n}=-1\}$. Consequently, for such $x$ we have $\mathbb{P}_{x}(\tau>n)=p^{n}$, and thus
$\lambda_{a}(p)=p\text{ for }a\in(0,\tfrac{1}{2}].$ (19)
This allows us to obtain the following analogue of Theorem 1 and Corollary 2 for all $x\geq 0$.
Proposition 4.
Let $\{X_{n}\}$ be a Markov chain defined by equation (1) with some $a\in(0,\frac{1}{2}]$. Assume that the innovations $\{\xi_{n}\}$ satisfy (4) with some $p\in(0,1)$. Denote
$\sigma^{\prime}=\inf\{n\geq 0:X_{n}<1/a\},\qquad\sigma^{\prime\prime}=\inf\{n\geq 0:X_{n}<6\}.$
Then for any $x\geq 0$ when $a<\frac{1}{2}$ and any $x\in[0,2)$ when $a=\frac{1}{2}$, for all integer $n$ large enough we have
$\mathbb{P}_{x}(\tau>n)=\mathbb{E}_{x}p^{-\sigma^{\prime}}\cdot p^{n},$ (20)
and for any $x\geq 2$ when $a=\frac{1}{2}$, we have
$\mathbb{P}_{x}(\tau>n)\sim q\mathbb{E}_{x}p^{-\sigma^{\prime\prime}}\cdot np^{n-1}$ (21)
as $n\to\infty$. Moreover, for any $x\geq 0$, the conditional distributions $\mathbb{P}_{x}(X_{n}\in\cdot\mid\tau>n)$ converge weakly to the $\delta$-measure at point $\frac{1}{1-a}$. This measure is quasi-stationary when $a<\frac{1}{2}$, in the sense of (18). There is no quasi-stationary probability measure when $a=\frac{1}{2}$.
Notably, the order of asymptotics in (21) differs from that in (14) and (20).
1.3. Outline of the approach and generalizations
Our study of the asymptotics of the persistence probability $\mathbb{P}_{x}(\tau>n)$ uses one of the most standard approaches via (quasi-) compactness of the transition operator $P$ of $\{X_{n}\}$ killed at leaving $0,\infty)$. The main novelty consists in the choice of an appropriate Banach space. Our choice is prompted by the connection between our AR(1)(1) chain with the $\pm 1$ innovations and the dynamical system given by the iterations of $T_{a}$. This is explained in detail in Section 2.3. We will see that $P$ is quasi-compact on a certain closed subspace of the Banach space $BV$ of functions of bounded variation on $0,\frac{1}{1-a}]$, where $P$ can be represented by a linear operator acting on summable sequences of length $\varkappa_{a}^{\prime}+1$; see Proposition 14 and Remark 15. We will use this representation to show that the leading eigenvalue of $P$ is $\lambda_{a}$ and the corresponding eigenfunction is $V$, and then prove (14). To prove the convergence of conditional distributions in (17), we consider the operator $P$ on the whole of $BV$ and derive an appropriate version of the Perron–Frobenius theorem for $P$, see Subsection 6.1. We will also show that $P$ is quasi-compact on $BV$; see Remark 20.
In all of our proofs it takes much more effort to consider the case $p<1/2$, where we need to use uniform upper bounds for the frequencies $L_{n}(x)/n$ of zeros in the $(\frac{1}{a},\frac{1}{2})$-expansion of $ax/2$. We believe that such bounds, presented in Section 3, are of independent interest.
Remark 5.
@@ 1.3. Outline of the approach and generalizations
Our approach can be extended to study persistence of the AR(1) chains with the innovations that take two arbitrary values of different sign. Indeed, thanks to a multiplicative rescaling, it suffices to consider the case where $\mathbb{P}(\xi_{1}=1)=p$ and $\mathbb{P}(\xi_{1}=-b)=q=1-p$ for some $b>0$. Then for any $a\in(\frac{b}{b+1},\frac{b+1}{b+2})$, the deterministic relation in reversed time (6) remains valid if we substitute $T_{a}$ by the mapping $T_{a,b}$ that is defined by
@@
Define $\varkappa_{a,b}(x)$ as in (7) with $I_{a}$ replaced by $(\frac{a(b+1)-b}{1-a},1)$.
It is easy to check that, for any $a$ as above, our proofs of (14) and (17) carry over without change, and these relations remain true when $\lambda_{a}$, $\delta_{k}$, etc. are replaced by the corresponding quantities $\lambda_{a,b}$, $\delta_{k}^{(b)}$, etc., defined in terms of $T_{a,b}$ and $\varkappa_{a,b}$ in place of $T_{a}$ and $\varkappa_{a}$. In general, it can be that $\varkappa_{a,b}=1$, as opposed to $\varkappa_{a}\geq 2$. Furthermore, it follows that $1-\delta_{k}^{(b)}(x)$, for $0\leq k<\varkappa_{a,b}(x)+1$, are the first digits in the $(\beta,\alpha)$-expansion of $\frac{ax}{b+1}$ with $\beta=\frac{1}{a}$ and $\alpha=\frac{b}{b+1}$. For the critical value $a=\frac{b+1}{b+2}$, our proofs of (14) and (17) remain valid without change if the orbit of 0 under $T_{a,b}$ does not hit $1$, yielding
$\lambda_{\frac{b+1}{b+2},b}\Big(\frac{1}{2}\Big)=\frac{b+2}{2(b+1)}.$
The structure of the rest of the paper is clear from the titles of the following sections.
2. Deterministic dynamics under time reversal and three approaches to persistence
In this section we prove the deterministic dynamics in the reverse time, given by (6), and describe the three possible approaches to persistence of the AR($1$) chains with $\pm 1$ innovations. All these approaches use the dynamical system featuring in (6). The first method is based on a reduction to finite Markov chains. Unfortunately, this reduction does not work for all values of the parameter $a$. For this reason we introduce an alternative, more analytic approach based on compactness properties of the operator $P$. The third approach works only for $a=\frac{2}{3}$. Although we do not make use of it, we present it to highlight a connection to the well-developed theory of transfer operators.
@@ 2. Deterministic dynamics under time reversal and three approaches to persistence
2.1. Deterministic behaviour in reverse time
@@
Let us prove (6) for $a\in[\frac{1}{2},\frac{2}{3})$. If $X_{n}<1$, then $X_{n}=aX_{n-1}-1$ because the other option $X_{n}=aX_{n-1}+1$ is impossible by $aX_{n-1}+1\geq 1$ on $\{\tau>n\}$. Hence $X_{n}\leq\frac{2a-1}{1-a}$ by (22). Thus, $X_{n}$ is in the domain of $T_{a}$ and we can write $X_{n-1}=\frac{1}{a}(X_{n}+1)=T_{a}(X_{n})$. If $X_{n}\geq 1$, then $X_{n}=aX_{n-1}+1$ because otherwise $X_{n}=aX_{n-1}-1<1$ by (22). Then $X_{n-1}=\frac{1}{a}(X_{n}-1)=T_{a}(X_{n})$. We thus checked that $X_{n-1}=T_{a}(X_{n})$ holds true on $\{\tau>n\}$ in either case, and (6) follows by induction.
@@
$\xi_{n-k}=(-1)^{1-d_{k}(aX_{n}/2)}=2d_{k}(aX_{n}/2)-1.$ (23)
Thus, $(\xi_{n-k}+1)/2$ for $0\leq k\leq n-1$ are the first $n$ digits in the $(\frac{1}{a},\frac{1}{2})$-expansion of $aX_{n}/2$.
The deterministic dynamics in the reversed time described above is essentially the reason for singular continuity of the quasi-stationary distributions $\nu_{a}$. The same effect can be also observed in the unconditional setting. More precisely, we shall now show that the unconditioned chain $\{X_{n}\}$ is deterministic in the reversed time for every $a\in(0,\frac{1}{2})$. We first notice that $|X_{k}|\geq\frac{1-2a}{1-a}$ for all $k\geq 1$ and for all starting points $x\in[-\frac{1}{1-a},\frac{1}{1-a}]$; this follows as in (22). Assume that $X_{n}\geq\frac{1-2a}{1-a}$. In general, we have two possibilities for $X_{n-1}$, namely $X_{n-1}=\frac{1}{a}(X_{n}+1)$ and $X_{n-1}=\frac{1}{a}(X_{n}-1)$. The assumption $X_{n}\geq\frac{1-2a}{1-a}$ implies that the case where $X_{n-1}=\frac{1}{a}(X_{n}+1)$ is not possible since
@@
This property implies that the limiting distribution $\pi_{a}$ is supported on the points $z$ whose orbit under $G_{a}$ does not hit the interval $\left(-\frac{1-2a}{1-a},\frac{1-2a}{1-a}\right)$. The set of such points has Lebesgue measure zero. Since the two “branches” of $G_{a}$ are onto the whole of the interval $\frac{1}{1-a},\frac{1}{1-a}]$, besides showing that the distribution $\pi_{a}$ is singular and continuous, one can compute the Hausdorff dimension of its support. The proof of these properties via deterministic evolution in the reversed time can be found in Lemma 2 by Bovier [9]. Relation (II-17) there is equivalent to (24).
2.2. Reduction to a finite chain when $\varkappa^{\prime}_{a}<\infty$